OpenStudy (anonymous):

How do you change the fraction 4x^3-8x^2+2x-1/2x into a mixed number?

OpenStudy (pfenn1):

Is this your expression?\[4x^3-8x^2+2x-\frac12x\]

OpenStudy (anonymous):

no it's 4x^3−8x2^+2x-1 over 2x

OpenStudy (pfenn1):

The whole expression over 2x like this?\[\frac{4x^3-8x^2+2x-1}{2x}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

how would you change that to a mixed expression?

OpenStudy (pfenn1):

this will equal\[\frac{4x^3}{2x}-\frac{8x^2}{2x}+\frac{2x}{2x}-\frac{1}{2x}\]

OpenStudy (anonymous):

i see! thanks!!!

OpenStudy (pfenn1):

you're welcome